### v1

# Hypergames: Modeling Misaligned Perceptions and Nested Beliefs for Multi-agent Systems

Abstract
Classical game-theoretic models typically assume rational agents, complete information, and common knowledge of payoffs – assumptions that are often violated in real-world multi-agent systems (MAS) characterized by uncertainty, misaligned perceptions, and nested beliefs. To overcome these limitations, researchers have proposed extensions that incorporate models of cognitive constraints, subjective beliefs, and heterogeneous reasoning. Among these, hypergame theory extends the classical paradigm by explicitly modeling agents’ subjective perceptions of the strategic scenario, known as perceptual games, in which agents may hold divergent beliefs about the structure, payoffs, or available actions. We present a systematic review of agent-compatible applications of hypergame theory, examining how its descriptive capabilities have been adapted to dynamic and interactive MAS contexts. We analyze 44 selected studies from cybersecurity, robotics, social simulation, communications, and general game-theoretic modeling. Building on a formal introduction to hypergame theory and its two major extensions – hierarchical hypergames and the hypergame normal form (HNF) – we develop agent-compatibility criteria and an agent-based classification framework to assess integration patterns and practical applicability. Our analysis reveals prevailing tendencies, including the prevalence of hierarchical and graph-based models in deceptive reasoning and the simplification of extensive theoretical frameworks in practical applications. We identify structural gaps, including the limited adoption of HNF-based models, the lack of formal hypergame languages, and unexplored opportunities for modeling human–agent and agent–agent misalignment. By synthesizing trends, challenges, and open research directions, this review provides a new roadmap for applying hypergame theory to enhance the realism and effectiveness of strategic modeling in dynamic multi-agent environments.
keywords:
Theory of Mind, Multi-Agent Systems, Game Theory, Hypergame Theory
1 Introduction
Strategic decision-making is a key issue in many real-world scenarios involving interactions among multiple agents, and as such is a fundamental agentic capability. Since agents in these settings must choose actions based on the actions of others, game theory naturally emerges as the discipline providing formal tools to model, design, and analyze such systems [58]. However, real-world environments are often marked by uncertainty, misaligned perceptions, and nested beliefs, which frequently render traditional models inadequate. To address these limitations, researchers have developed frameworks that incorporate cognitive constraints, recursive reasoning, and heterogeneous agents exhibiting bounded rationality [33, 63, 54, 15, 20].
Simpler and less expressive models, such as the k-level theory and the cognitive hierarchy model [15], acknowledge differences in players’ reasoning capacities but offer limited insight into the structure and underlying nature of these differences. Epistemic game theory provides a formal language for representing player types, belief hierarchies, and uncertainty; however, its predominantly probabilistic models only partially account for the mechanisms that give rise to these cognitive variations [20].
In contrast, hypergame theory advances the game-theoretic paradigm by explicitly modeling subjective representations of the strategic scenario – so-called perceptual games – in which agents may disagree about the components of the game: the participating players, the available actions, or the payoff structure [7]. This capability makes hypergames particularly suitable for multi-agent system (MAS) frameworks that aim to emulate human-like reasoning and cognitive sophistication. Such suitability aligns with recent interest in theory of mind (ToM) – a psychological model of human reasoning that characterizes the ability to attribute mental states to oneself and others [57] – and its integration into agentic artificial intelligence [59].
Originally developed as an analytical framework for post-hoc conflict analysis, hypergame theory was not designed for dynamic or interactive agent-based systems. As a result, despite its descriptive power and emerging practical applications, its full potential in MAS environments remains underutilized. This systematic review aims to address that gap by identifying existing agent-compatible applications of hypergame theory, analyzing patterns of practical integration, and highlighting both structural challenges and emerging opportunities at the intersection of hypergames and multi-agent systems.
We make several contributions. First, we provide a formal introduction to hypergame theory, including its foundational concepts and its two major revisions: hierarchical hypergames and the hypergame normal form (HNF). Building on this foundation, we define agent-compatibility criteria for hypergame models and apply these criteria in a systematic review of 44 selected papers from the hypergame-theoretic literature. We develop an agent-based classification framework to examine how these works integrate hypergame-theoretic models, categorizing them by domain and sub-domain, practical applicability, and degree of integration – from conceptual inspiration through partial implementation to full deployment. We also construct a structured taxonomy of hypergame formalisms, distinguishing multi-level and HNF models, and map their usage across planning, reasoning, learning, and uncertainty modeling tasks. Furthermore, we synthesize trends and usage patterns, showing that while theoretical integrations of hypergames remain prevalent, practical deployments often rely on simplified or flattened models. Finally, we discuss gaps in infrastructure, such as the lack of agent-oriented hypergame modeling languages, and suggest avenues to address these gaps by drawing inspiration from formal languages like epistemic logic and GDL-III, which are already used in MAS research.
The remainder of this paper is organized as follows. Section 2 reviews the foundations of hypergame theory, including its game-theoretic roots, the formalization of games, and the development of hierarchical and HNF-based hypergame models. Section 3 presents the methodology and results of our systematic review, covering publication trends, a taxonomy of hypergame models, and an analysis of their application in practice – including application domains, integration methodologies, fidelity, and task alignment. Section 4 discusses the survey results, with particular attention to usage patterns, domain-model associations, and the need for agent-based hypergame languages. Section 5 situates this work within the broader literature, and Section 6 concludes with a summary of findings and directions for future research.
2 Hypergame Theory
In this section, we first establish the foundations of game-theoretic reasoning by introducing classical models of strategic interaction. We then introduce the theory of hypergames as an extension of game theory that relaxes the assumptions of common knowledge and shared perceptions. Finally, we present two major formalisms of hypergame theory: the Hypergame Normal Form (HNF) and Hierarchical Hypergames, both of which have served as theoretical bases for modeling misaligned perceptions and recursive reasoning in multi-agent systems. For the purpose of introducing the relevant concepts, we use the two-player rock-paper-scissors game as a running example – a well-known, simple setting where both players must simultaneously choose one of three available actions: Rock, Paper, or Scissors.
2.1 Game-theoretic Foundations
While decision theory provides tools to reason about how an individual agent can derive choices, when another decision-maker is involved the player may have to explicitly reason about how the opponent is trying to solve the same problem. Game theory is better equipped for such analyses, as it deals with conflicts where decisions and their outcomes are interdependent between agents [53].
2.1.1 Elements of a Game
[58] introduces the key elements of game-theoretic interactions through the acronym PAPI: Players, Actions, Payoff and Information. We define these four elements as follows:
Players
Players are the decision-making entities involved in the strategic interaction. Game-theoretic models typically assume that players are [53]:
  • Rational: they consistently act to maximize their own utility functions;
  • Strategically intelligent: they understand the structure of the game and anticipate that others behave similarly.
Actions
Actions represent the immediate choices available to each player in the game they are involved with. In contrast, strategies are agents’ plans of actions based on the observed interaction history [56]. We distinguish between two classes of strategies [51]:
  • Pure: deterministic mappings from states or histories to actions;
  • Mixed: probability distributions over pure strategies.
Payoffs
Payoffs quantify the preferences of each player over possible outcomes. They are typically represented as real-valued utility functions $U_{i}:A\to\mathbb{R}$, where $A$ is the joint action space. Numerical payoffs enable the use of optimization and equilibrium analysis [78, 30].
Information
Information concerns players’ knowledge at the time of making a decision. This knowledge encompasses a player’s image of the other players, the options available to everyone, and the historical moves. We classify games based on three information criteria:
  • Perfect vs. Imperfect Information: Whether agents can fully observe others’ actions [56];
  • Complete vs. Incomplete Information: Whether agents can fully observe others’ payoffs [30];
  • Symmetric vs. Asymmetric Information: Whether agents can access the same quality and quantity of information [58].
2.1.2 Formalizing Games
A game-theoretic model of rock-paper-scissors can be formally defined as a tuple $G=(N,A,u)$ where:
  • $N=\{1,2\}$ is the set of players
  • $A=A_{1}\times A_{2}$ is the set of joint actions, where for each player $i\in N$:
    • $A_{i}=\{\text{Rock},\text{Paper},\text{Scissors}\}$ is the action set of player $i$
    • A pure strategy for player $i$ is an element $a_{i}\in A_{i}$
    • A mixed strategy for player $i$ is a probability distribution $\sigma_{i}$ over $A_{i}$
  • $u=(u_{1},u_{2})$ are the utility functions where for each player $i\in N$:
    $u_{i}:A\rightarrow\mathbb{R}\text{ defined as: }$:
    $u_{i}(a_{i},a_{j})=\begin{cases}1&\text{if }(a_{i},a_{j})\in W_{i}\\ 0&\text{if }a_{i}=a_{j}\\ -1&\text{if }(a_{i},a_{j})\in L_{i}\end{cases}$
    where:
    • $W_{i}=\{(\text{Rock},\text{Scissors}),(\text{Paper},\text{Rock}),(\text{Scissors},\text{Paper})\}$ are winning outcomes
    • $L_{i}=\{(\text{Rock},\text{Paper}),(\text{Paper},\text{Scissors}),(\text{Scissors},\text{Rock})\}$ are losing outcomes
Expected Utility
When players employ mixed strategies – probability distributions over their pure actions – the outcome of the game becomes probabilistic. The expected utility for a player is then calculated as the average payoff, weighted by the likelihood of each joint action occurring under the players’ chosen strategies [78]. Formally, given mixed strategies $\sigma_{i}$ and $\sigma_{j}$ for players $i$ and $j$, the expected utility for player $i$ is:
$EU_{i}(\sigma_{i},\sigma_{j})=\sum_{a_{i}\in A_{i}}\sum_{a_{j}\in A_{j}}\sigma_{i}(a_{i})\sigma_{j}(a_{j})u_{i}(a_{i},a_{j})$
Rational players are typically assumed to choose strategies that maximize their expected utility, forming the basis of solution concepts such as the mixed-strategy Nash equilibrium.
Nash Equilibrium
[…]
$u_{i}(a_{i}^{*},a_{-i}^{*})\geq u_{i}(a_{i},a_{-i}^{*})\quad\forall a_{i}\in A_{i}$
In case of mixed strategies, the Mixed Strategy Nash Eqilibrium (MSNE) is defined analogously: each player’s mixed strategy must be a best response to the mixed strategies of others. Formally:
$EU_{i}(\sigma^{*}_{i},\sigma^{*}_{-i})\geq EU_{i}(\sigma_{i},\sigma^{*}_{-i})\quad\forall\sigma_{i}\in\Delta(A_{i}),$
where $\Delta(A_{i})$ denotes the set of mixed strategies over $A_{i}$.
In the case of rock-paper-scissors, no pure strategy profile constitutes a Nash equilibrium. However, a unique mixed strategy Nash equilibrium exists where each player chooses Rock, Paper, and Scissors with equal probability [10].
2.2 Games of Misperception
Real-world agents may differ in their abilities to observe and process information, potentially leading to distinct internal models of others and misaligned understandings of the conflict, the available actions, or the associated outcomes. Hypergames provide an explicit model of such perceptions by capturing each agent’s subjective interpretation of the situation [8]. Furthermore, hypergame analyses rely on preference relations rather than numerical payoffs, resulting in a qualitative model of desirability. This approach aims to mitigate challenges associated with assigning precise numerical values to real-world outcomes.
Hypergames are considered models of bounded rationality [63, 76] because they support reasoning in which an agent may perceive an opponent’s choice as irrational, even though that choice is rational from the opponent’s own perspective [45]. Formally, a hypergame is composed of each player’s perceptual game – a version of the conflict that represents the player’s interpretation – and is defined as the tuple $H=(N,\{G_{i}\}_{i\in N})$ where:
  • $N=\{1,2,\ldots,n\}$ is the set of agents;
  • For each agent $i\in N$, $G_{i}$ represents agent $i$’s perceived game, where $G_{i}=(N_{i},A_{i},R_{i})$ with:
    • $N_{i}\subseteq N$ is the set of agents as perceived by $i$;
    • $A_{i}=\bigtimes_{j\in N_{i}}A_{ij}$ is the joint action space as perceived by $i$, where:
        $A_{ij}$ is $i$’s perception of $j$’s available actions;
    • $R_{i}=\{R_{ij}\}_{j\in N_{i}}$ is the set of preference relations as perceived by $i$, where:
        $R_{ij}\subseteq A_{i}\times A_{j}$ is $i$’s perception of $j$’s preference relation;
        For outcomes $x,y\in A_{i}$, $x>y$ indicates that $i$ prefers outcome $x$ to outcome $y$.
Note: elements with indices $ii$ denote player $i^{\prime}s$ own interpretation of the game’s components.
To illustrate misaligned perceptions, consider an asymmetric version of the rock-paper-scissors game modeled as a hypergame in which one agent’s action space is constrained. Let $H=(N,\{G_{1},G_{2}\})$, where:
  • $G_{2}$ is Agent 1’s perceived game:
    • $N_{1}=\{1,2\}$;
    • $A_{11}=A_{12}=\{\text{Rock},\text{Paper},\text{Scissors}\}$;
    • $R_{11}=R_{12}=$
      $\{(P,R),(S,P),(R,S)\}>\{(R,R),(P,P),(S,S)\}>\{(S,R),(R,P),(P,S)\};$
  • $G_{2}$ is Agent 2’s perceived game:
    • $N_{2}=\{1,2\}$;
    • $A_{21}=A_{22}=\{\text{Rock},\text{Paper}\}$;
    • $R_{21}=R_{22}=$
      $\{(R,R),(P,P)\}>\{(R,P),(P,R)\};$
This example captures a situation where Scissors is not recognized as an available action by Agent 2, whereas Agent 1 perceives the complete action space. This form of perceptual asymmetry mirrors the classic ”Fall of France” hypergame analysis [6]. The resulting discrepancy in perceived game structure can lead to unexpected outcomes – referred to in hypergame theory as strategic surprise [8].
2.3 Hierarchical Hypergames
Hypergame theory introduces a framework for modeling individual agents’ subjective interpretations of the conflict they are involved in. By incorporating higher-order perceptions [8], hypergames enable the formalization of recursive reasoning – e.g., a decision-maker $r$ might consider what $q$ believes about how $p$ perceives the situation. These are called $m$-th order expectations, where $m$ corresponds to the depth of the perceptual chain.  [81] later formalized and extended Bennet’s framework into a hierarchical hypergame analysis structure, addressing the overly abstract nature of earlier definitions and proposing a more rigorous, standardized representation of multi-level perceptual games. In this extended formalism, a $0^{th}$-level hypergame is equivalent to a standard game-theoretic without misaligned perceptions and is defined as a tuple $H^{0}=(N,P,S,U,V)$, where:
  • $N=\{i\mid 1\leq i\leq n\}$ is the set of players;
  • $P=\{P_{i}\mid i\in N\}$ is the set of players option sets, where:
    • $P_{i}$ is the set of actions available to player $i$;
  • $S=\{S_{i}\mid i\in N\}$ is the set of players’ strategy sets, where:
    • $S_{i}\subseteq P_{i}$ is the set of actions player $i$ considers for $G$;
  • $U$ is the set of all possible outcomes, where:
    • Each outcome $u\in U$ is a tuple of chosen strategies $u=(s_{1},\ldots,s_{n})$;
    • $U=\bigtimes_{i\in N}S_{i}$ represents all possible strategy combinations;
  • $V=\{V_{i}\mid i\in N\}$ is the set of player preferences, where for each player $i$:
    • $V_{i}=(U,R_{i})$ where $R_{i}$ is a preference relation over $U$, or
    • $V_{i}=(U,u_{i})$ where $u_{i}:U\rightarrow\mathbb{R}$ is a utility function.
The distinction between $P_{i}$ and $S_{i}$ allows the modeling of agents’ beliefs about physical or cognitive constraints. A zero-level hypergame assumes shared perceptions – there are no subjective differences between agents’ views of the game.
Higher-order Reasoning
To support higher-order reasoning, Wang et al. introduced the notion of perceptual functions. Each player’s viewpoint encodes their holistic interpretation of the game, while individual perceptions capture their interpretation of specific components. For any player $i$, a perception is defined as a mapping:
$f_{i}:\Gamma_{i}\to\Gamma_{ij},$
where:
  • $\Gamma_{j}$ is a component from player $j$’s game (from $\{N,P,S,U,V\}$);
  • $\Gamma_{ij}$ denotes player $i$’s internal model of $\Gamma_{j}$;
  • For each element $\gamma\in\Gamma_{j}$, $\varphi=f_{i}(\gamma)$ is $i$’s perceived image of $\gamma$.
In the case of nested beliefs, player $i$’s interpretation is composed via the product mapping:
$f_{i}\circ f_{j}:\Gamma_{k}\rightarrow\Gamma_{ijk}.$
The resulting image is the product of individual perceptual functions:
$\varphi=f_{i}(f_{j}(\gamma))=f_{i}\circ f_{j}(\gamma)=f_{ij}(\gamma).$
Misperceptions can arise across all game components when the image $\varphi=f_{i}(\gamma)$ deviates from the true $\gamma\in\Gamma_{j}$. A perception is:
  • Underperceived if it omits elements present in $\Gamma_{j}$;
  • Overperceived if it includes elements not present in $\Gamma_{j}$;
  • Otherwise, a misperception if $\varphi\neq\gamma$.
Higher-order Games
In cases where players develop misaligned perceptions and expectations about others’ games, we define higher-order hypergames as follows:
  • A first-level hypergame assumes that players are unaware of any perceptual misalignment. Each player $i$’ viewpoint is captured by their own perceptual game $G_{i}$, forming the composite hypergame $H^{1}=\{G_{i}\mid 1\leq i\leq n\}$. Each perceptual game is defined using potentially distinct sets $N_{i},P_{i},S_{i},U_{i},V_{i}$.
  • A second-level hypergame arises when at least one player is aware that others may have different viewpoints. $H^{2}=\{H_{i}^{1}\mid 1\leq i\leq n\}$, where each $H^{1}_{i}$ now contains the set of perceptual games $G_{ij}=(N_{ij},P_{ij},S_{ij},U_{ij},V_{ij})$ that player $i$ believes player $j$ uses: $H_{i}^{1}=\{G_{ij}\mid 1\leq j\leq m\}$.
  • A third-level hypergame extends this recursion once more: a player $i$ models what player $j$ believes about what player $k$’ game. We have $H^{3}=\{H_{i}^{2}\mid 1\leq i\leq n\}$, where each individual game is composed of first-level hypergames $H_{i}^{2}=\{H_{ij}^{1}\mid 1\leq j\leq m\}$ and each first-level hypergame is a set of standard games $H_{ij}^{1}=\{G_{ijk}\mid 1\leq k\leq h\}$, where each percpetual game is $G_{ijk}=(N_{ijk},P_{ijk},S_{ijk},U_{ijk},V_{ijk})$.
At each level, the formal structure is extended by an additional layer of subjective (hyper)games built on nested perceptions. While the mathematical framework supports $L^{th}$-order hypergames of arbitrary depth, we restrict the discussion to level 3, reflecting empirical observations that human strategic reasoning seldom exceeds three levels [15] and balancing expressiveness and computational tractability.
Hypergame Nash Equilibrium
The Hypergame Nash Equilibrium (HNE) extends the classical Nash Equilibrium to settings where players may have divergent perceptions of the game [81]. The HNE formalizes mutual best responses in this context of misaligned perceptions; that is, the HNE is an NE in each player’s corresponding subjective game. Given a strategy profile $a^{*}=(a^{*}_{i},a^{*}_{-i})$:
  • In a first-level hypergame $H^{1}=\{G_{i}\mid 1\leq i\leq n\}$, $a^{*}$ is a HNE if, for each player $i\in N$, $a^{*}$ is a NE in their subjective game $G_{i}$;
  • In a second-level hypergame $H^{2}=\{H^{1}_{i}\mid 1\leq i\leq n\}$, $a^{*}$ is a HNE if, for each player $i\in N$ and $j\in N_{i}$, $a^{*}$ is a NE in $G_{ij}$;
  • In a third-level hypergame $H^{3}=\{H_{i}^{2}\mid 1\leq i\leq n\}$, $a^{*}$ is a HNE if, for each player $i\in N$, $j\in N_{i}$ and $k\in N_{ij}$, $a^{*}$ is a NE in $G_{ijk}$.
2.4 Hypergame Normal Form
The normal form is a foundational representation in game theory used to express strategic interactions among decision-makers, where players’ actions and the corresponding payoffs are captured in a table, typically referred to as the payoff matrix [58]. The Hypergame Normal Form (HNF) extends the standard normal form to capture a single player’s beliefs – drawing on decision-theoretic concepts [32] – about their opponent’s reasoning [77]. Originally proposed for analyzing complex military planning scenarios, HNF enables an agent to reason over multiple possible opponent strategies, each corresponding to a different perceptual subgame. Table 1 illustrates an HNF matrix for the rock-paper-scissors game used in previous examples from the row player’s perspective. The HNF consists of:
  • The Full Game (bottom right quadrant) represents the “expert’s view” – a standard $n\times m$ game-theoretic payoff matrix, where $n$ and $m$ denote the set of actions available to the row and column players, respectively;
  • The Column Mixed Strategies (top right quadrant) or CMS represent possible opponent strategies, including:
    • $\mathbf{C_{0}}$: the Nash equilibrium mixed strategy (NEMS);
    • $\mathbf{(C_{1},\ldots,C_{k})}$: the row player’s belief-induced subgames representing alternative opponent mixed strategies over $(S_{1},\ldots,S_{m})$;
    • $\mathbf{C_{\Sigma}}$: a weighted summary of beliefs, computed as:
      $S_{j}=\sum_{k=0}^{K-1}P_{k}c_{kj},\quad\text{for }j=0,1,\ldots,n$
      where $P_{k}$ is the weight for belief context $k$ and $c_{kj}$ is the weight for strategy $j$ in context $k$.
  • The Row’s Belief Context (upper left quadrant) contains probabilities assigned to each opponent CMS. These are analogous to perceptual functions in multi-level hypergames but focused on strategy selection rather than game structure. They capture the player’s beliefs about the likelihood of each opponent’s reasoning mode, that is, the probability that the corresponding subgame (CMS row) is active;
  • The Row Mixed Strategy (RMS) zone (bottom left quadrant) captures the row player’s responses to each belief context, resembling second-order beliefs in multi-level hypergames. RMS entries are not prescriptive but inform planning.
Column Mixed Strategy
Row’s Belief Context $C_{\Sigma}$ $S_{1}$ $S_{2}$ $S_{3}$
$P_{2}$ 0.5 $C_{2}$ $\frac{1}{2}$ $\frac{1}{2}$
$P_{1}$ 0.25 $C_{1}$ 0.326 0.3456 0.3284
$P_{0}$ 0.25 $C_{0}$ $\frac{1}{3}$ $\frac{1}{3}$ $\frac{1}{3}$
$H_{1}$ $R_{2}$ $R_{1}$ $R_{0}$ Rock Paper Scissors
$H_{11}$ 0.2 0.3284 $\frac{1}{3}$ Rock 0 -1 1
$H_{12}$ 0.6 0.326 $\frac{1}{3}$ Paper 1 0 -1
$H_{13}$ 0.2 0.3456 $\frac{1}{3}$ Scissors -1 1 0
Row Mixed Strategy Full Game
Table 1: Hypergame Normal Form matrix for rock-paper-scissors comprising of: (1) Row’s Belief Context with probabilities $P_{k}$ for each belief context, (2) Column Mixed Strategies showing opponent mixed strategies $C_{k}$ and their aggregate $C_{\Sigma}$, (3) Row Mixed Strategy containing $R_{k}$ responses to each belief context, and (4) Full Game with cardinal utilities. Hyperstrategies $H_{ij}$ represent evaluated plans, expressed as probability distributions over row strategies. Each $C_{k}$ and $R_{k}$ represents a mixed strategy vector summing to 1.
The game in Table 1 defines the row player’s beliefs about the column player in rock-paper-scissors. $C_{0}$ corresponds to the mixed strategy Nash equilibrium (uniform randomization), while $C_{1}$ reflects empirical distributions derived from online gameplay data [10]. In $C_{2}$, the row player believes that the opponent cannot play Scissors. The belief context assigns the highest probability to $C_{2}$, indicating that the row player expects this restricted subgame to be most likely. The RMS entries mirror responses to each $C_{k}$. Additionally, the HNF matrix can include expected utility (EU) and hypergame expected utility (HEU) calculations as additional rows beneath the matrix – omitted in Table 1 for clarity.
Hyperstrategies
Hyperstrategies are mixed strategies over RMSs, represented as vectors of assigned probabilities adjacent to the RMS quadrant. We include a hypothetical hyperstrategy $H_{1}$ in Table 1 for illustration. Vane [77] defines four hyperstrategies that meet the following efficiency criterion:
$\text{Eff}(hs)\leftrightarrow\text{EU}(hs,C_{\Sigma})\geq\text{EU}(\text{NEMS}(m\times n)).$
These are defined as follows:
  • NEMS: NEMS is the baseline Nash equilibrium mixed strategy derived from the full game;
  • Modeling Opponent (MO): MO is a decision-theoretic strategy selecting the best response based on CMS and RMS utilities;
  • Pick Subgame (PS): PS selects the Nash equilibrium of the subgame identified by MO;
  • Weighted Subgame (WS): WS is a hybrid strategy that chooses the best option based on the combination of the MO-specified CMS and $R_{0}$ weighted by the corresponding belief context probabilities.
Hypergame Expected Utility
Vane [75] introduced hypergame expected utility (HEU) as a bridge between decision theory and game theory. HEU incorporates a parameter $g$ – the “fear of being outguessed” – and the worst-case outcome $G$, defined as the lowest expected utility for the row player. Then, given a hyperstrategy $hs$, we define the HEU as:
$\text{HEU}(hs,g)=(1-g)\cdot\text{EU}(hs,C_{\Sigma})+g\cdot\text{EU}(hs,G).$
The relative performance of hyperstrategies such as MO, PS, and WS varies with the uncertainty parameter $g$, with NEMS becoming increasingly dominant as uncertainty rises. The analysis of $g$’s effects – detailed in [75] – suggests that when agents hold accurate beliefs about their opponent’s strategy, HEU-optimized hyperstrategies can outperform standard equilibrium analysis.
3 Hypergame Theory in Multi-Agent Systems
While hypergame theory is a model of bounded rationality that can provide multi-agent systems with a framework that can model agent heterogeneity and environmental uncertainties to a greater level of detail than probabilistic methods, it was originally designed to be an analytical framework targeting post-hoc studies of complex conflicts instead of providing tools for the decision-maker on the spot. While hypergame models have yet to become prevalent in dynamic settings multi-agent systems entail, they have a potential to overcome the constraints posed by typical game-theoretic models on the complexity agent-based frameworks can capture [61]. Motivated by this identified potential, in this section we provide a systematic review of hypergame-theoretic works focusing on studies that entail agent-based applications or the potential to be integrated into dynamic agentic systems.
3.1 Methodology
Our systematic survey of hypergame theory in the context of multi-agent systems follows a multi-step screening and classification process. Our goal was to identify works where hypergame-theoretic models are not only referenced, but explicitly applied in ways that support agentic reasoning, decision-making, or strategic interaction within MAS-compatible contexts. Figure 1 provides an overview of this selection funnel. Starting with an inclusive set of keyword-matched results, we progressively filtered out duplicates, non-English works, irrelevant uses, and purely analytical studies lacking actionable agent-level integration. The steps are outlined below:
Initial Collection
We conducted a keyword-based query based on Google Scholar using the term “hypergame theory”, which yielded 320 unique results as of June 2025.
Duplicate Removal
We removed 303030 entries due to duplicity. Removed entries included different versions of the same manuscript (e.g., preprints and published versions) as well as overlapping content such as thesis chapters and standalone publications.
Language Filter
We excluded 17 papers for being written in languages other than English.
Survey and Meta-reviews
We identified 6 publications that were either surveys of hypergame theory or broader reviews that discussed hypergames as part of a more general framework. These were retained for comparative purposes but not included in the core set analyzed for model classification.
Relevance Filtering
While hypergame theory is a distinct subfield within game theory, the term is occasionally used in unrelated contexts or cited without substantive application. To ensure relevance, we manually examined each entry’s abstract, methods, references, and in-text keyword usage. We excluded 154 papers that either lacked an actual implementation of hypergames, used the concept only metaphorically, or mentioned it in passing without integration into the modeling framework.
Agent-compatibility Filter
From the remaining 113 works, we applied a final inclusion criterion focused on agent compatibility. Specifically, we retained only those works in which hypergames were used to support or influence agent behavior, strategic decision-making, or reasoning under uncertainty – either in simulated, practical, or theoretical settings with agent-level implications. This step excluded 69 additional papers that employed hypergames solely from an analyst perspective, or in a way that did not feed back into the agent’s own actions or perceptions.
As a result, we identified a final corpus of 44 papers featuring agent-compatible hypergame-theoretic models. Although only three of these explicitly formalize the agent architecture, the majority assume rational decision-makers with embedded utilities or subjective games. Due to the inconsistency in architectural specification, the focus of our review excludes implicit and explicit agent design.
Figure 1: Distribution of the evaluated 320 papers based on our selection criteria. The left chart illustrates the composition of papers eliminated due to irrelevance, duplication, or being written in a language other than English. The chart to the right illustrates the agent-compatibility evaluation that resulted in the final 44 papers considered for our review.
3.2 Hypergame-theoretic Publication Trends
Bennett introduced the theory of hypergames as an analytical framework extending standard game-theoretic frameworks, emphasizing the modeling ability to account for perceptual misalignments. Vane’s hypergame normal form revised Bennett’s hypergame representation, providing an extended decision-theoretic matrix form focusing on one player’s belief systems and reasoning. Early applications of both frameworks solely entailed post-hoc or hypothetical analyses, with no demonstrations on how a decision-maker can benefit from hypergame theory in real-time. While both original authors repeatedly highlighted the potential of hypergames to be used in dynamic agent-based applications and both made attempts of “live” hypergame analyses, the adoption of hypergame theory, in multi-agent literature in particular, remains to take off. Nonetheless, our review reveals a trend of growing number of MAS-related hypergame-theoretic publications in recent years. We see a shift towards agent-based applications, as illustrated on Figure 2.
Figure 2: Frequency of all hypergame-relevant and agentic publications over time.
Interestingly, [45] stated that “Unfortunately very little in the way of hypergame theory application has been done in the arena of cyber warfare”, while a decade later the current shift is dominantly lead by attack-defense modeling and simulations in cybersecurity – as demonstrated by Figures 679 and the various cybersecurity-related surveys covering hypergame-based approaches we discuss in section 5.
Hypergame theory is a niche field that has not yet gained traction, with a modest number of publications, dominated by a historical concentration of contributions from the original authors and their collaborators. To investigate whether recent trends – such as increasing domain diversity and methodological variation – are accompanied by a broader dissemination of authorship, we conducted an analysis of author frequencies and co-authorship patterns across the 44 papers reviewed. We extracted all named authors from each publication, accounting for different name formats and removing placeholder or malformed entries. First, we identified individual authors associated with at least three of the reviewed papers.
Figure 3: Publication counts per authors with at least 3 associated works.
Then, to identify collaborative groups, we constructed a co-authorship graph in which nodes represent individual authors and edges indicate co-authorship on at least one paper. Groups were then defined using connected components of this graph, with each component representing a set of authors linked by direct or indirect collaboration. For each group, we computed the number of papers coauthored by at least two members, thereby capturing the extent and productivity of distinct author clusters.
Table 2: Identified coauthor groups with at least 3 shared papers. Member names are shown by surname only.
Group Shared Papers Members
G5 11
Cohen, Fried, Fu, Gaglione, Han, Hemida, Kamhoua,
Kulkarni, Leslie, Li, Liu, Luo, Ma, Shi, Udupa, Xi
G6 6 Cortés, Gharesifard
G1 4
Anwar, Cho, Kamhoua, Singh, Wan, Zhu
G8 3
Chen, Cheng, Gao, Guan, He, Hong, Ma, Yuan
As shown in Figure 3, ten authors are associated with at least three papers, with nearly a third of reviewed papers attributable to two recurring author pairs: Kulkarni & Fu and Gharesifard & Cortés. Furthermore, as captured by Table 2, we found four groups of collaborators with at least three relevant publications, where Group 5 – Kulkarni & Fu et al. – is responsible for a quarter of publications. At the same time, over half of the papers originate from authors or groups with fewer than three publications, suggesting a shift from concentrated research programs toward more diverse, application-specific uses of hypergames, potentially reflecting growing interest and wider adoption of the theory across domains – a finding well-aligned with the publication trends identified via our classification framework and Figure 2.
3.3 Taxonomy of Hypergame Models
In this section, we provide a detailed taxonomy of hypergame models identified across the surveyed papers. We categorize the models according to three key dimensions: the foundational hypergame concept and game dynamics, formalizations of hypergame frameworks, and the computational tasks hypergame-theoretic approaches address. This taxonomy facilitates a structured understanding of how different hypergame implementations vary in their theoretical assumptions, representational complexity, and practical application scenarios, enabling researchers to select or develop hypergame models tailored to specific multi-agent system requirements.
Game Dynamics
First, we looked at whether the hypergame models entail a single round of interaction or repeated, iterative decision-making. The distribution across hypergame concepts and game dynamics is characterized by 30 applications that entail repeated interactions, while 14 models rely on one-shot games. This suggests a shift in applications trends, as historically works primarily involved one-shot post-hoc studies.
Hypergame Concept
We also distinguished works by whether they were inspired by the original [8] or extended [81] hypergame framework – collectively denoted as Multi-level Hypergames (MLH) – or based on the Hypergame Normal Form (HNF). We found 7 papers that refer to HNF, 35 rely on Multi-level Hypergames, and 2 papers use components from both concepts, which we denote as Hybrid approaches.
Hypergame Formalizations
Next, we identified common practical hypergame model definitions, which we classify as follows:
  • HNF-based solutions populate complete Hypergame Normal Form (HNF) matrices, which agents directly utilize for decision-making.
  • HEU-based methods utilize the Hypergame Expected Utility derived from HNF to help agents select optimal strategies.
  • L-th Order Multi-level Hypergames populate hierarchical hypergame structures, enabling agents to form nested beliefs.
  • Flattened L-th Order Multi-level Hypergames employ subjective games but exclude the hierarchical tree structure.
  • Bayesian L-th Order Hypergames implement hybrid belief models in which hypergames represent perceptions, but agents perform decision-making using Bayesian reasoning.
  • Perceptual Game-based implementations enable agents to develop hypergame-like beliefs without defining a complete multi-level hypergame structure.
  • HNE-based frameworks involve decision-makers explicitly relying on the Hypergame Nash Equilibrium concept.
  • Graph-based methods incorporate spatial hypergames and strategies utilizing graph-based representations.
  • Stackelberg Hypergames extend the leader-follower mechanism of Stackelberg games by integrating hypergame-theoretic subjective perceptions.
The distribution of model usage across the surveyed papers is depicted in Figure 4: the majority of works implement a model based on the typical multi-level hypergame concept (L-th Order, Flattened L-th Order and Bayesian L-th Order) or hypergames on graphs. 5 works entail hybrid solution, where in addition to populating a hypergame model a hypergame-theoretic solution concept (HEU or HNE) is also implemented. In two cases, the hybrid approach entailed cross-framework implementations: multi-level-hypergame-based models integrating the hypergame expected utility concept from HNF.
Figure 4: Distribution of model types. Hybrid models that span multiple categories are counted in each relevant type.
3.4 Hypergames in Practice
Hypergames, despite being originally developed as a post-hoc analytical framework, provide a promising tool for modeling realistic interaction scenarios with misaligned perceptions in multi-agent systems. Practical implementations employing, to varying extents, hypergame concepts emerge in such diverse fields as cybersecurity, robotics, and social simulations. To identify current patterns in the practical implementation of hypergames, as well as the challenges and open research problems, we analyze the papers that implement hypergame concepts in terms of the application domain (Section 3.4.1), hypergame integration methodology (Section 3.4.2), integration fidelity (Section 3.4.3), and finally, the role of the hypergame model within the given framework (Section 3.4.4).
3.4.1 Application Domains
To investigate how hypergame theory has been adopted across different areas of research and application, we classified the 44 agent-compatible papers according to their primary domain of use. Each paper was assigned to one of six high-level domains, which reflect the thematic focus of the study: Game Theory, Cybersecurity, Social Simulations, Robotics, Multi-agent Systems, and Communications. Within each domain, we identified recurring sub-domains that capture specific use cases, each briefly introduced in the following subsections and highlighted in bold to aid navigation. In Figure 5, we provide an overview of the reviewed works across the identified domains and sub-domains, which are summarized as follows.
Figure 5: Distribution of agent-compatible hypergame papers across high-level application domains and their associated sub-domains, where sum-domains can overlap between application domains.
Game Theory Works that advance game theory in general, without applying to any specific practical domain, are categorized under the general domain of Game Theory. These papers utilize hypergame-theoretic tools to define equilibrium concepts, model learning dynamics under misperception, or devise strategic responses to asymmetry and deception. For example,  [64] integrates Fuzzy Game Theory with hypergames by incorporating fuzzy strategies and imprecise preference perceptions. In contrast,  [83] focuses on Deception and uses hypergames to analyze how misinformation affects equilibrium strategies. Additionally,  [52] proposes a hypergame-based Q-learning approach to solve Pursuit-Evasion games, while  [62] develops a hypergame-theoretic model of deception in concurrent Stochastic Games. Lastly, various works leverage the hypergame-theoretic Evolution of Beliefs in adversarial games characterized by misperceptions. These studies define mechanisms for belief updates [24, 28] that target the learning of equilibria [25, 26] and belief manipulation [27].
Cybersecurity Cybersecurity represents the most active domain for agent-compatible hypergame applications. This prevalence can be attributed to the inherent suitability of hypergames for capturing misinformation and modeling active deception – both critical issues within cybersecurity contexts. The majority of the surveyed works address Attack-Defense Modeling in general, formulating hypergames on graphs and applying linear temporal logic to guide deceptive strategy synthesis [41, 42, 47, 50]. Other examples include belief manipulation [29, 4], mission impact assessments employing hypergame-theoretic frameworks [73, 84], and other specialized hypergame-based methodologies for attack-defense modeling [34, 23, 5, 14]. In addition, several works target a narrower area of attack-defense modeling. Defensive Deception entails the use of deceptive strategies to mislead the attacker, often focusing on optimizing strategic honeypot, decoy, and false target assignment [82, 43, 79, 2, 39, 66]. Alternative approaches include using defensive deception strategies to increase the cost of attacks through modeling uncertainty via hypergames [37] and maximizing strategic surprise via sensor revealing strategies [74], or exploring two-way misinformation in deception games [17] and defensive deception against multiple attackers [80].
Social Simulations Social simulations involve multi-agent-based modeling of strategic interactions within conceptualized social environments. The framework proposed by [69] employs a hypergame-based Theory of Mind reasoning framework. This approach has been utilized by [70] to explore the ability of agentic large language models (LLMs) to engage in human-like reasoning during guessing games. Furthermore, [46] leverages hypergame theory to model Coalition Detection in natural language reasoning tasks performed by LLMs.
Robotics In the realm of robotic Control Theory, hypergame applications, for example, focus on modeling obstructed perceptions in traffic scenarios using subjective games [48]. In contrast, [40] aim to develop hypergame-based strategies in asymmetrical information environments between robots and their surroundings. Other studies in robotics explore the implementation of hypergames in the context of Autonomous Vehicles, particularly in the design of deceptive defense strategies for unmanned ground vehicles that operate under asymmetric perception conditions [49]. Others address the cognitive alignment of unmanned aerial vehicles during defensive [35, 36] or exploratory [18] missions.
Multi-agent Systems We categorize applications related to general issues within MAS under this domain. [1] defines a set of hidden-role games targeting deceptive objectives and equipping agents with second-order Theory of Mind conceptualized as second-order hypergames. This framework allows agents to act based on misaligned perceptions of their roles and objectives, facilitating the emergence of asymmetric beliefs and belief manipulation within partially observable, mixed cooperative-competitive settings.
Communications Within the communications sector, we examine decentralized multi-user wireless networks composed of transmitters and receivers whose interactions are limited by bandwidth, channel quality, and computational resources. Semantic communication described a paradigm where only the most task-relevant information is transmitted while receivers utilize reasoning and reconstruction techniques to infer any missing data. In this context, [71] applies hypergame theory to model and manage perceptual discrepancies among agents explicitly.
3.4.2 Hypergame Integration Methodology
During the paper selection process, we have already selected a subset of papers that satisfy the agent compatibility criteria. We then further classified these works based on the practicality of their proposed implementation, as follows.
  • Theoretical works that extend or use hypergame theory in an application domain relying on agency. While these works do not entail a runnable system or dynamic experiments, an applicable algorithmic description and/or a practical case is provided.
  • Experimental works entail hypergame-based applications that implement an experimental setting and use empirical evidence to prove or evaluate the effects of hypergame-based approaches.
  • Practical applications encapsulate works that implement hypergame frameworks that are deployable and are tested in realistic simulations and real-life inspired settings.
Figure 6 reveals that even in our surveyed set of works that we filtered based on agent-compatibility theoretical papers still contribute to half of all relevant publications, while experimental and practical works share a count of 11 each. Even across the more practical domains – cybersecurity, robotics, social simulations – theoretical papers prevail. This significant focus on the theoretical application of hypergames is well-aligned with the original design of hypergame theory as an analytical framework and with the challenges we articulated previously.
Figure 6: Domain-specific distribution of agent-compatible hypergame papers by application type. Theoretical works are conceptually grounded and contain algorithmic descriptions or formalized case examples, but lack implementation or simulation. Experimental works implement hypergame-based models in controlled settings to evaluate their effects empirically. Practical applications deploy fully integrated hypergame frameworks in realistic simulations or real-world-inspired scenarios.
3.4.3 Hypergame Integration Fidelity
Next, we distinguish works based on the extent of their hypergame integration as follows:
  • Conceptual applications refer to hypergames and are inspired by hypergame-theoretic concepts but do not incorporate an explicit hypergame model.
  • Partial applications are inspired by hypergame-theoretic concepts but only formally integrate components – such as solution concepts or subjective perceptions – instead of instantiating the entire framework;
  • Complete applications entail a full hypergame model integration.
As depicted by Figure 7, we can observe that 35 of surveyed papers entail a complete integration of hypergame-theoretic models, with 5 works relying on partial and 3 on conceptual implementations. Furthermore, the “complete” category is dominant across individual domains, suggesting that while the operationalization of the complex hypergame analytical frameworks is non-trivial, agent-based application benefit from high-fidelity instantiations of hypergame-theoretic models.
Figure 7: Domain-specific distribution of agent-compatible hypergame papers based on the extent of model integration. Conceptual applications reference hypergame-theoretic ideas but do not instantiate explicit models. Partial applications formally incorporate some hypergame components such as subjective perceptions or equilibrium concepts, without realizing a complete framework. Complete applications implement fully specified hypergame models within the agentic context.
3.4.4 Hypergame Task
Finally, we classify hypergame models based on their role in the framework; what is the task the approach is seeking to solve via hypergame theory? The classification entails planning, uncertainty measure, learning, and reasoning. Models addressing multiple tasks were counted for each class. The overall distribution – as shown in Figure 8 – depicts that the majority of works utilize hypergame-theoretic models for reasoning – or decision-making in general – or complementing learning algorithms.
Figure 8: Distribution of surveyed papers by primary computational task implemented by hypergame-theoretic models.
4 Discussion of Survey Results
Table 3: Summary of agent-compatible hypergame paper classification based on model taxonomy and model application details.
Hypergame Model Taxonomy Hypergame Application Details
Paper
Dynamics
Concept
Formalization
Domain [^1] CS: Cybersecurity, SS: Social Simulations, GT: Game Theory, COM: Communications, ROB: Robotics
Sub-
domain [^2] DD: Defensive Deception, ADM: Attack-Defense Modeling, TOM: Thoery of Mind, CD: Coalition Detection, EOB: Evolution of Beliefs, P-E: Pursuit-Evasion, SC: Semantic Communication, AV: Autonomous Vehicles, CT: Control Theory
Method
Fidelity
Task
[37] Repeated HNF HEU-based CS DD Practical Partial
Reasoning,
Uncertainty
Measure
[23] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[82] Repeated MLH L-th Order CS DD Experimental Complete Reasoning
[72] Repeated MLH Stackelberg CS ADM Experimental Complete Learning
[42] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[69] One-shot MLH
Flattened
L-th Order
SS TOM Practical Complete Reasoning
[47] Repeated MLH Graph-based CS ADM Theoretical Complete Reasoning
[43] One-shot MLH Graph-based CS DD Theoretical Complete Reasoning
[46] One-shot MLH L-th Order SS CD Theoretical Complete Learning
[50] Repeated MLH Graph-based SS ADM Theoretical Complete Planning
[83] Repeated MLH HNE-based GT DC Theoretical Conceptual Reasoning
[28] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[24] Repeated MLH Graph-based GT EOB Theoretical Complete Learning
[25] Repeated MLH Graph-based GT EOB Theoretical Complete
Learning,
Uncertainty
Measure
[79] Repeated HNF
Stackelberg,
HEU-based
CS DD Practical Partial
Reasoning,
Uncertainty
Measure
[2] Repeated HNF
HEU-based
CS DD Practical Partial
Reasoning
[52] Repeated MLH
Flattened,
L-th Order
GT P-E Experimental Complete
Reasoning,
Learning
[34] Repeated HNF HNF CS ADM Experimental Complete Learning
[71] Repeated MLH Stackelberg COM SC Experimental Complete Learning
[18] Repeated MLH Perceptual ROB AV Practical Conceptual Planning
[48] One-shot MLH L-th Order ROB CT Practical Complete
Uncertainty
Measure
[39] Repeated MLH Graph-based CS DD Experimental Complete Reasoning
[66] Repeated
MLH,
HNF
L-th Order,
HEU-based
CS DD Practical Complete Reasoning
[73] Repeated HNF HEU-based CS ADM Experimental Partial
Reasoning,
Learning
[35] One-shot MLH Stackelberg ROB AV Theoretical Complete Reasoning
[…]
Table 4: (Cont.) Summary of agent-compatible hypergame paper classification based on model taxonomy and model application details.
Hypergame Model Taxonomy Hypergame Application Details
Paper
Dynamics
Concept
Formalization
Domain [^1] CS: Cybersecurity, GT: Game Theory, MAS: Multi-agent Systems, ROB: Robotics, SS: Social Simulations
Sub-
domain [^2] ADM: Attack-Defense Modeling, EOB: Evolution of Beliefs, TOM: Thoery of Mind, DD: Defensive Deception, CT: Control theory, AV: Autonomous Vehicles, TD: Trojan Detection, FUZ: Fuzzy Game Theory
Method
Fidelity
Task
[4] Repeated MLH
L-th Order,
HNE-based
CS ADM Theoretical Complete Reasoning
[26] Repeated MLH Graph-based GT EOB Theoretical Complete
Reasoning,
Learning
[1] Repeated MLH
Bayesian
L-th Order
MAS TOM Practical Conceptual
Reasoning,
Learning
[5] One-shot MLH L-th Order CS ADM Theoretical Complete Reasoning
[17] Repeated HNF HEU-based CS DD Experimental Complete
Reasoning,
Uncertainty
Measure
[40] One-shot MLH L-th Order ROB CT Theoretical Complete
Reasoning,
Planning
[49] Repeated MLH Graph-based ROB AV Practical Complete Reasoning
[62] Repeated MLH L-th Order GT SG Experimental Complete Planning
[74] Repeated MLH Graph-based CS DD Theoretical Complete Reasoning
[80] Repeated
MLH,
HNF
L-th Order,
HEU-based
CS DD Practical Complete
Reasoning,
Learning
[14] One-shot MLH
Stackelberg,
HNE-based
CS ADM Theoretical Complete Reasoning
[29] Repeated MLH Graph-based CS ADM Theoretical Complete
Reasoning,
Learning
[27] Repeated MLH Graph-based GT EOB Theoretical Complete
Reasoning,
Learning
[84] One-shot HNF HEU-based CS ADM Practical Partial
Reasoning,
Uncertainty
Measure
[41] One-shot MLH Graph-based CS ADM Theoretical Complete
Reasoning
[70] One-shot MLH
Flattened
L-th Order
SS TOM Experimental Conceptual
Reasoning
[36] One-shot MLH Stackelberg ROB AV Theoretical Complete
Reasoning
[19] Repeated MLH L-th Order CS TD Experimental Complete
Reasoning
[64] One-shot MLH L-th Order GT FUZ Theoretical Complete
Reasoning
[…]
In the preceding sections, we established a comprehensive taxonomy of hypergame models and explored various applications that utilize these models. The full classification of the 44 reviewed papers is documented in Tables 3 and 4. In this section, we reflect on emerging patterns and structural gaps identified through our systematic review of hypergame-theoretic literature and the classifications disclosed in the previous sections, aiming to contextualize the current state of the field and identify fruitful directions for future research.
4.1 Hypergame Usage Patterns
In this section, we analyze emerging patterns in the design and usage of hypergame models across agentic applications. Drawing on the classification dimensions outlined earlier, we identify recurring choices in model structures, domain integration, and computational tasks. These patterns offer practical insights for researchers aiming to deploy hypergame-theoretic mechanisms in dynamic multi-agent systems.
4.1.1 Overview of Model Usage
An evaluation of the core concept (multi-level hypergame vs. HNF) usage across surveyed works at a domain-level decomposition (Figure 9) reveals that multi-level hypergames are the dominant framework across all categories, and only cybersecurity works entail implementations of HNF or hybrid models.
Figure 9: Distribution of core hypergame concepts implemented per application domain. HNF entails works inspired by Vane’s hypergame normal form [77], Multi-level Hypergames are based on Bennet’s original concept [8] or Wang et al.’s extended formalism [81]. Hybrid approaches use concepts from both main concepts.
At a lower-level view, graph-based hypergames and L-th order hierarchical hypergames dominate (Figure 4). Graph-based models in particular stand out as the most popular implementation, often supporting spatial planning and adversarial reasoning, especially in cybersecurity and general game-theoretic applications (Figure 10).
Figure 10: Distribution of hypergame model types across application domains. Hybrid models that span multiple categories are counted in each relevant type.
However, when theoretical applications are omitted, shifting the focus toward experimental and practical implementations that are closer to deployable multi-agent systems, the distribution changes considerably. As depicted in Figure 11, 8 out of the 14 works we classified as practical entail simplified hypergame-theoretic concepts – perceptual games, flattened L-th order hypergames or HEU – instead of integrating and populating models entirely according to the complete hypergame analyses frameworks. These findings suggest, that while practical, real-world applications can benefit from hypergame-theoretic models, their integration must entail concretizations of the primarily analytical frameworks. Figures 6 and 12 paint a similar picture, where practical application from domains more aligned with real-world applicability – cybersecurity, robotics, social simulations and general MAS – works seem to benefit most from the simplified models.
Figure 11: Distribution of experimental and practical hypergame applications across model types.
Figure 12: Distribution of hypergame model across application domains, omitting theoretical applications.
Meanwhile, HNF has seen limited use in agentic contexts. Only 9 works adopt HNF-based solutions (Figure 9), with all situated in the cybersecurity domain. This suggests that HNF’s fixed-matrix structure may be too rigid or low-level for modern agent-based modeling, where higher-order and domain-specific reasoning demands greater representational flexibility.
4.1.2 Task-Model Alignment
Figure 13: Distribution of model types per hypergame task.
As we can see on Figure 13, reasoning is the most frequent application, with 33 papers using hypergames to support decision-making under misaligned perceptions. Hypergame-based learning tasks are present in about a third of the surveyed papers, most frequently associated with graph-based and HEU-based models, which offer scalable formulations that lend themselves to integration with learning algorithms. Apart from one paper, estimating uncertainty appears as a secondary task, complementing reasoning or learning, with mostly probabilistic, HEU-based approaches. The limited presence of hypergames in high-level planning tasks may reflect the absence of expressive, agent-compatible formalisms akin to epistemic planning languages.
To complement these findings, we evaluate how different task-based works are composed of theoretical, experimental and practical model applications (Figure 14) and revisited the distribution of hypergame model types per tasks (Figures 612). While practical applications are the majority for deriving hypergame-based uncertainty measures, for planning, learning and reasoning-related tasks purely theoretical and more applicable works have an equal split.
Figure 14: Distribution of hypergame task with composition of theoretical, experimental and practical applications.
4.1.3 Domain-Model Associations
When decomposed by application domain (Figure 10), distinct preferences in model types emerge:
  • Cybersecurity leads in diversity and volume, encompassing both L-th order and HNF-based models, as well as graph-based and Stackelberg hypergames or hypergame equilibrium-based appraoches. The prevalence of deception-centric scenarios appears to benefit from nested and multi-layered perception modeling, making hypergames a natural fit.
  • General game-theoretic works entail mostly theoretical but high-fidelity, complete integrations of hypergame models.
  • Social simulations and multi-agent system studies predominantly use flattened L-th order and L-th order hypergames, highlighting a focus on recursive reasoning rather than spatial deception. This aligns with their frequent overlap with theory of mind literature and epistemic modeling.
  • Robotics and communications rely primarily on subjective perceptual representations, which can encode environmental uncertainties and cognitive bias, particularly in navigation and defense tasks.
4.2 Agent-based Hypergame Languages
While the extended formalisms by [81] and [77] concretized hypergame frameworks and provided guidelines for hypergame analysis, there is no unified formal language for reasoning about hypergames. Motivated by this gap, hypergame markup language (HML) was developed by [12] to formally encode hypergame models for his hypergame analysis software: HYPANT. HML supports subjective subgames based on player perception but is limited to static utilities and does not support functional utility definitions or dynamic belief updates. Another hypergame analysis software, HAT [31], relies on XML-based structured representation of hypergames. While HAT supports iterative runs and belief updates, it relies on populating Vane’s hypergame normal form representation-based structures. [45] also mention the Security Policy Assistant (SPA) developed by Vane to assist in sensitive decision-making, such as evaluating foreign disclosure of classified information. SPA is not publicly available and we could not find any material for a more detailed analysis.
Despite the aforementioned tools and preliminary modeling languages, an agent-based hypergame modeling language or simulation platform has not emerged. While we encountered various works in our survey that integrate hypergame-theoretic models with linear-temporal-logic-based planning, they do not define a hypergame-specific description language. Future hypergame modeling efforts may take inspiration from alternative frameworks that address information asymmetry and nested beliefs to develop a hypergame-theoretic formal language grounded in logic, such as epistemic logic model that enable dynamic belief updates and planning with iterated knowledge [13, 3], GDL-III providing a formal specification language for general game playing with games of incomplete information [67], or TOMA extending epistemic logic to formalize the theory of mind [21].
5 Related Work
While Bennett initially developed hypergame theory as an analytical tool for structuring complex strategic scenarios, subsequent research has explored more dynamic applications. [9] demonstrated the use of hypergame theory in real-time decision-making during a shipping crisis. Later, the development of Vane’s HNF provided a more methodological version with potential applications in dynamic, agent-based models. The potential is articulated in his paper addressing decision-theoretic agents [75] and his brief, high-level review of hypergames [76], where the connection to the Observe-Orient-Decide-Act paradigm was also established – a concept well-aligned with human-cognition-inspired agent theories such as the belief-desire-intention (BDI) architecture [68, 65]. Jointly, including the additional formalism provided by [81], while these advancements of hypergame theory offer a richly formal and generalizable framework, up to this point, this line of research remains unconnected to the agent-based and multi-agent systems literature.
[45] presented a broad review of hypergame theory as a framework for modeling conflict, misperception, and deception, particularly in adversarial domains like military planning and cyber defense. Their review emphasizes the theoretical basis of hypergames and highlights how layered beliefs and misaligned perceptions give rise to strategic opportunities unavailable in classical game theory or decision theory. However, their review primarily focused on illustrative, mostly analytical examples rather than methodological analyses.
[85] offered a dual-framework perspective, reviewing defensive deception from both machine learning and game-theoretic viewpoints. Although hypergame theory is featured in their survey through selected analytical and practical examples, the coverage remains limited and application-specific, without substantial linkage to broader multi-agent systems research. Complementing this work, [38] presented a systematic review of game-theoretic defense models against advanced persistent threats. While not dedicated to hypergames and, as such, lacking formal descriptions of hypergame models, their review acknowledges the value of hypergame theory in modeling adversarial perturbations and covers three additional hypergame-based works on defensive deception. Crook’s doctoral dissertation [16] represents a unique contribution bridging computability theory and hypergames, involving an exploration of how attack graphs in cybersecurity can be recast as hypergames. While preliminary, this work proposes several directions for formalizing adaptive, layered-belief hypergames in cybersecurity with links to computable analysis. Finally, [22] addressed hypergame theory in their review of game-theoretic models in cyber deception. While their discussion is brief and only on a high level, the paper explicitly recognizes that attacker–defender interactions often exhibit asymmetries traditional games cannot capture. The consideration of hypergames for the adaptability of defender strategies progresses the practical generalizability of hypergame theory.
To our knowledge, no prior work provides a systematic survey of hypergame theory from the perspective of multi-agent systems. Existing reviews typically focus on early decision-theoretic formalisms, defensive deception in cyber-physical systems, or specific ad-hoc and post-hoc case studies. Our survey addresses this gap by:
  • Providing a rigorous classification framework explicitly tailored to multi-agent systems, emphasizing works that implement and apply hypergame theory within agent-compatible contexts;
  • Presenting a structured taxonomy of hypergame models across multiple dimensions, including formal hypergame structures, agent architectures, domain-specific implementations, and computational tasks;
  • Identifying key patterns in hypergame integration, such as the prevalence of cybersecurity-related hypergame-theoretic applications, while also highlighting and discussing reasons for the limited use of frameworks like Hypergame Normal Form (HNF) in other domains;
  • Mapping the explicit relationship between hypergame model types and their deployment in reasoning, planning, learning, and uncertainty management tasks, thereby underscoring the alignment between hypergame formalisms and their intended agentic functionalities;
  • Demonstrating the evolving landscape of hypergame theory, shifting from historically concentrated theoretical analyses toward more diverse, practical deployments across multiple MAS-relevant domains;
  • Identifying structural gaps and open research challenges, particularly in hypergame-based epistemic modeling, human–AI interaction, and high-level planning frameworks.
By synthesizing these insights, our survey not only bridges existing knowledge gaps but also provides a clear roadmap for future hypergame-theoretic research, particularly within dynamic, agent-driven multi-agent systems. The outlined trends and classifications serve as actionable references, encouraging researchers to leverage hypergames in practical MAS applications and inspiring further methodological innovation within the hypergame theory community.
6 Conclusions
In this work, we conducted a systematic review of hypergame theory, focusing on its adaptation and application within multi-agent systems (MAS). We surveyed 44 papers spanning multiple domains and analyzed how hypergame models have been integrated into agent-based frameworks, examining trends, formalizations, and application patterns. Our review highlighted both the theoretical strengths of hypergame theory – particularly its ability to represent nested beliefs, misaligned perceptions, and subjective representations of strategic interactions – and the practical challenges of deploying such models in dynamic MAS environments.
Despite early efforts such as HML [12] and HAT [31], there remains no formal language or simulation framework explicitly designed for modeling agent-based hypergames. Given the identified lack of a unified representational formalism and the need for structured support for recursive belief modeling and dynamic perception updates, future work should prioritize the development of dedicated hypergame description languages tailored to agent-based applications. Inspiration may be drawn from planning frameworks rooted in epistemic logic [13, 3], general game-playing languages like GDL-III [67], or formal ToM modeling tools such as TOMA [21].
Across the surveyed domains, hypergames have been successfully employed to model nested beliefs, misaligned perceptions, and deception; however, few works explicitly link these models to established cognitive agent architectures. As recursive reasoning becomes increasingly central in agent design, integrating hypergame theory with human-inspired frameworks such as BDI [68, 65] could enrich symbolic reasoning and offer interpretable mechanisms for belief formation and adaptation in dynamic environments. Furthermore, while we encountered only a handful of hypergame-based approaches addressing large language models (LLMs), the growing prevalence of human–agent [60] and agent–agent misalignment [44], along with the application of theory of mind in agentic systems [11, 59, 21], suggests promising opportunities for hypergame theory to formalize ToM-like models capable of capturing and mitigating agent misalignment.
Finally, several surveyed works highlighted hybrid reasoning approaches, such as Bayesian $L$-th order hypergames, yet few explored the design space that combines subjective perception modeling with probabilistic inference. Current HNF-based approaches have not yet provided a general solution for dynamically populating the HNF matrix. Future research could explore hybrid agent architectures that integrate hypergame structures for belief representation with probabilistic reasoning modules for decision-making, enabling agents to better navigate uncertain, misperceived, or adversarial contexts.
We hope that this survey will guide and inspire future researchers seeking to model misperceptions, nested beliefs, and recursive reasoning in multi-agent-based applications. By identifying trends, gaps, and trajectories in the literature, we aim to catalyze the development of hypergame-based AI models that are both theoretically grounded and practically applicable.
7 Declarations
[…]
Conflict of interest The authors declare no conflict of interest.
[…]
Funding This work was supported by a Leverhulme Trust International Professorship Grant (LIP-2022-001).
[…]
Acknowledgement The authors would like to express their gratitude to David Levine for his constructive comments. The first author would also like to thank him for his supervision and the fruitful discussion of game-theoretic and hypergame-theoretic materials.
[…]
Open Access This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons licence, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons licence, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons licence and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this licence, visit http://creativecommons.org/licenses/by/4.0/.